OpenStudy (anonymous):

find f(a+h) - f(a) ___________ h for f(x) = 1 _____ x-3

OpenStudy (anonymous):

you are basically plugging in F(x) everywhere you see a.

OpenStudy (anonymous):

actually that is not quite correct

OpenStudy (anonymous):

\[f(a)=\frac{1}{a-3}\] \[f(a+h)=\frac{1}{a+h-3}\] \[f(a+h)-f(a)=\frac{1}{a+h-3}-\frac{1}{a-3}\] \[=\frac{-h}{(a+h-3)(a-3)}\]

OpenStudy (anonymous):

never seen it done that way, but ok I trust you. Thanks

OpenStudy (anonymous):

then divide by h and get \[\frac{-1}{(a+h-3)(a-3)}\]

OpenStudy (anonymous):

the weird thing is none of my possible answers have a -1..but all have 1s

OpenStudy (anonymous):

hold on let me check my algebra

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

answer i wrote it correct. you are probably getting ready to take the derivative, which would be doing the same algebra and then letting h go to zero, and you should get \[-\frac{1}{(a-3)^2}\]so i know it is correct. perhaps your possible answer is \[-\frac{1}{(a+h-3)(a-3)}\] with the minus sign out front instead of next to the 1 in the numerator it is the same thing

OpenStudy (anonymous):

THANK YOU SO MUCH!!!